# Instruction Fusion: Advancing Prompt Evolution through Hybridization

### 25 Dec 2023 v1

@@ 2.2 Instruction Fusion
Fusion Process. Figure 1 gives an illustration of the Instruction Fusion process. Denote the initial dataset as $\mathbf{S}=\{(I_{i},R_{i})\}_{1<i<N}$, where $I_{i}$, $R_{i}$ are the $i-{th}$ seed instruction and the corresponding response, $N$ is the number of instructions in $\mathbf{S}$. To create a new instruction using Instruction Fusion, we first random select two seed instructions $i_{j}$ and $i_{k}$ from $I$, where $1\leq j,k\leq N$. Then, we prompt GPT4-Turbo to fuse the selected two instructions into $I^{m}_{(j,k)}$ utilizing the following prompt:
@@
For each fused instruction, if $I^{m}_{(j,k)}==\text{INVALID\ PROMPT}$, we will discard the $(j,k)$ combination and randomly sample a new one. We keep repeating the fusion process untill the number of new instructions reach the target amount $M$. The result instructions set is then $\textbf{C}^{\text{IF}}=\{(I^{m}_{(j,k)},R^{m}_{(j,k)})\}_{1<m<M}$, where $R^{m}_{(j,k)}$ is the corresponding response of $I^{m}_{(j,k)}$ generated by GPT4-Turbo.
[…]
@@ 2.3 Complement the Deficiencies
Figure 3 also exemplifies how Instruction Fusion (IF) smoothing the difficulty gradient between original seeds and fused instructions. In the figure, consider a seed instruction that asks to print "Hello, World!"; Evol-Instruct evolves this by adding a filtering task, indeed increasing difficulty and diversity. However, this can lead to a difficulty gap, as the new requirement might not be present in other seeds. With IF, a combined instruction involving both filtering and printing only arises if both objectives are present in the seed instructions, ensuring a smoother difficulty gradient.
The better gradient can also be proved by the theory in (Kung et al., 2023)’ works. They categorizes samples with high instruction uncertainty and a high prediction probability as “ambiguous” samples. Instruction uncertainty refers to the extent to which minor changes in the instruction can influence the response generation. Greater uncertainty implies that LLMs are more sensitive to the sample’s instruction. In their study, it is highlighted that “ambiguous” training data, to which LLMs are particularly responsive, play a crucial role during fine-tuning. In contrast, samples with low instruction uncertainty and low prediction probability represent difficult samples, which could be to hard for LLMs to learn and has fewer benefits for fine-tuning.
[…]
@@
Instruction Fusion (IF) counters the diversity limitations of Evol-Instruct, which is constrained by initial objectives, by creating instructions that blend the targets of parent instructions. This leads to unique tasks not found in the Evol-Instruct dataset. In Figure 5, the blue points represent a t-SNE 2D sentiment embedding plot for PI instructions, showing small cliques indicative of limited diversity. Conversely, the red points, representing IF’s fused instructions, populate previously empty areas, demonstrating increased diversity. The red points’ distribution fills gaps left by the blue points, illustrating a more uniform sentiment embedding. To quantify this dispersion, we calculate the variance of nearest neighbor distances for all instructions:
@@
The results in Figure 5 reveal notable findings. The first plot of $\textbf{C}_{PI}$, composed of $\textbf{C}_{30}+\textbf{C}_{30r}$, exhibits higher dispersion, indicating reduced uniformity compared to $\textbf{C}_{30}$ alone. However, an interesting contrast arises when employing Instruction Fusion on $\textbf{C}_{30}$: by integrating $\textbf{H}_{20}$ with $\textbf{C}_{30}$, the dispersion of the second plot decreases by 28.5%. This significant reduction suggests that Instruction Fusion effectively narrows the gaps in sentiment embeddings between existing instructions. This improvement is achieved by combining the objectives of two distinct instructions, demonstrating the method’s efficacy in enhancing instruction diversity.
In conclusion, the proposed Instruction Fusion (IF) method effectively complements the limitations of Evol-Instruct by addressing its shortcomings by integrating diverse objectives from different seed instructions, leading to a more challenging and varied instruction set. This approach not only broadens the scope of tasks but also ensures a more balanced and coherent progression in instruction complexity.
@@ 3 Experiments
IF-CL-EC 34B weight&data 75.0 70.1 73.7 63.9
In this section, we report the experiment details of LLMs fine-tuned with instrcution generated by IF. We focused on the pass@1 performance under greedy generation settings across five benchmarks: four Python benchmarks (HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), ${}^{+}$, ${}^{+}$ Liu et al. (2023a)) and the multi-language benchmark MultiPL-E Cassano et al. (2022). The official EvalPlus code [^3] https://github.com/evalplus/evalplus was used for evaluating the Python benchmarks, while bigcode-evaluation-harness Ben Allal et al. (2022) was utilized to evaluate the performance on MultiPL-E.
@@ 3.2 Evaluation
Table 2 presents the performance of various open-source models on Python benchmarks. When compared to the leading closed-source models, GPT-3.5 Turbo, our models with 13B and 34B parameters demonstrate superior performance on the HumanEval and ${}^{+}$ benchmarks. However, there is a noticeable performance gap on the two MPBB benchmarks. This discrepancy may stem from the MBPP test cases, which often require specific prior knowledge for resolution. For instance, in the real test case task_id:20, the prompt "Write a function to check if a given number is woodball or not." requires the model’s understanding of what constitutes a ’woodball’ number. This type of prior knowledge is either not acquired during the pre-training stage or forgotten during the supervised fine-tuning process, leading to a consistent failure across all Instruction Fusion (IF) models for this case.
However, compared with all open-source models, the Instruction Fusion models excel in Python code generation performance, surpassing all baselines. The fused instructions, which offer richer sentiment, increased difficulty, and smoother transitions, allow both the 13B and 34B versions of IF models to achieve new benchmarks, outperforming models of comparable sizes. Notably, the 13B version of IF-CLP-ECP outperforms all 34B baselines in the HumanEval and ${}^{+}$ benchmarks and shows comparable results on MBPP and ${}^{+}$.
@@ 3.3 Ablation Study
Method HumanEval HumanEval+ MBPP MBPP+ IF-CLP-ECP 73.8 69.5 71.7 61.7   -fused inst. 67.7 64.0 66.4 56.4 IF-CL-ECP 75.6 68.9 67.9 58.6   -fused inst. 66.5 61.6 67.2 57.4 IF-CL-EC 74.4 68.3 69.7 59.4   -fused inst. 65.2 61.0 68.4 56.1 IF-CLP-ECP 75.6 69.5 73.7 62.7   -fused inst. 72.0 65.2 72.2 61.9 IF-CL-ECP 76.8 70.7 73.2 62.9   -fused inst. 67.7 62.8 70.9 61.2 IF-CL-EC 75.0 70.1 73.7 63.9   -fused inst. 67.7 62.2 69.9 60.4
@@
Further analysis of Instruction Fusion’s effectiveness involved an ablation study using various combinations of IF and the original evolved instructions. Figure 6 depicts the performance of Code Llama when fine-tuned with different volumes of original instructions. The data indicates that the 13B model achieves optimal performance with 30K evolution samples, suggesting that an increased quantity of evolved instructions does not linearly translate to better performance. Moreover, Luo et al. (2023) observed a performance decrease when extending evolution to the fourth round, implying a tangible upper limit to the effectiveness of Code Evol-Instruct in code generation tasks. However, the integration of IF samples with the original evolved dataset seems to transcend this limitation, resulting in a notable enhancement in model performance. Conversely, the 34B model does not show the same constraints with the current range of evolved samples, possibly due to its larger parameter size.
@@ 5 Conclusion
We introduce the Instruction Fusion technique to address the limitations of Evol-Instruct in code generation tasks. This method prompts the teacher LLM (GPT-4 Turbo) to merge two evolved instructions into a unified, integrated prompt. This fusion process combines two distinct objectives into a single, more complex goal, which 1) increases the task’s difficulty in a reasonable manner by prompting Code LLMs to simultaneously achieve two diverse objectives, 2) enhances the variety within the instruction pool, and 3) creates a smoother difficulty gradient, as the objectives can be learned independently from the initial seed instruction. Experimental results demonstrate that LLMs fine-tuned with IF data excel in the five most commonly used benchmarks, setting new state-of-the-art performance standards. Furthermore, this method proves to be a valuable complement to Evol-Instruct.
@@
The primary limitation of the Instruction Fusion method proposed in our study is its cost. We employ GPT-4 Turbo as the Teacher Language Model (LLM), and the fusion prompt incorporates a higher token count compared to the Evol-Instruction. This results in increased expenses for data collection. However, it is worth noting compared with manual annotation. Moreover, such costs are rapidly decreasing, thanks to the swift advancements in LLM technology. Another point of concern is the success rate of the fusion process. Fusion of Python-based instructions achieves a pass rate of 93% (where 7% of the fused instructions are deemed unsolvable by GPT-4 Turbo). In contrast, cross-language fusion exhibits a lower pass rate of approximately 63%. This discrepancy arises because many tasks are typically resolved using specific programming languages. Consequently, fusing instructions from different programming languages often leads to impractical outcomes. To address this issue, we propose the development of new prompts that categorize tasks prior to sampling, a solution which could be to explore in future research.

### 27 Dec 2023 v2

@@ 2.2 Instruction Fusion
Fusion Process. Figure 1 gives an illustration of the Instruction Fusion process. Denote the initial dataset as $\mathbf{C}=\{(I_{i},R_{i})\}_{1<i<N}$, where $I_{i}$, $R_{i}$ are the $i^{\text{th}}$ seed instruction and the corresponding response, $N$ is the number of instructions in $\mathbf{C}$. To create a new instruction using Instruction Fusion, we first random select two seed instructions $I_{j}$ and $I_{k}$ from $\mathbf{C}$, where $1\leq j,k\leq N$. Then, denote the target amount of fused instructions as $M$, we prompt GPT-4 Turbo to fuse the selected two instructions into the $m^{\text{th}}$ fused instruction $I^{m}_{(j,k)}$ utilizing the following prompt:
@@
Here, $1\leq m\leq M$. For each fused instruction, if $I^{m}_{(j,k)}==\text{INVALID\ PROMPT}$, we will discard the $(j,k)$ combination and randomly sample a new one. We keep repeating the fusion process until the number of new instructions reach the target amount $M$. The result instructions set is then $\textbf{H}^{\text{IF}}=\{(I^{m}_{(j,k)},R^{m}_{(j,k)})\}_{1<m<M}$, where $R^{m}_{(j,k)}$ is the corresponding response of $I^{m}_{(j,k)}$ generated by GPT-4 Turbo.
Data Collection. Table 1 gives the statistic of the collected datasets. Due to the unavailability of WizardCoder’s Luo et al. (2023) original evolved instructions, we utilized the third-party dataset evol-codealpaca-v1111https://github.com/theblackcat102/evol-dataset, which mirrors the methods used by Luo et al. (2023). This dataset, comprising 111k refined instructions, is evolved from the CodeAlpaca [^2] https://github.com/sahil280114/codealpaca dataset, which is also serves as the seeds in WizardCoder. We divided evol-codealpaca-v1 based on programming language into Python-related instructions $\textbf{C}_{\text{PI}}$ and Non-Python-related instructions $\textbf{C}_{\text{NPI}}$, numbering 50,131 and 61,140, respectively. To perform detailed comparison in later experiments, we further divide $\textbf{C}_{\text{PI}}$ into $\textbf{C}_{30}$ and $\textbf{C}_{30r}$, which represents the 30K random samples from $\textbf{C}_{\text{PI}}$ and the rest samples, respectively. Based on the Instruction Fusion scheme introduced above, we set $\textbf{C}_{\text{PI}}$ as seeds and perform IF with $M=50,000$ to obtain the result dataset $\textbf{H}_{\text{PI}}$, where 20K fused instructions $\textbf{H}_{20}$ only use seeds from $\textbf{C}_{30}$. Then we perform another IF with $M=50,000$ on $\textbf{C}_{\text{PI}}+\textbf{C}_{\text{NPI}}$ and obtain $\textbf{H}_{\text{PI}+\text{NPI}}$. For cost efficiency, our method predominantly utilizes samples in $\textbf{H}_{\text{PI}}$. Nevertheless, towards the experiment’s end, we exam code generation performance using both two datasets, showcasing our method’s capability across multiple programming languages.
@@ 2.3 Complement the Deficiencies
Figure 3 illustrates the role of Instruction Fusion (IF) in mitigating the difficulty gradient between original seeds and fused instructions. For example, consider a basic seed instruction to print “Hello, World!”. Evol-Instruct evolves this by incorporating a filtering task, thereby elevating the difficulty and diversity. However, this evolution may introduce a disparity in difficulty, as the new element might be absent in other seeds. IF addresses this by ensuring new sub-objectives are introduced only if they exist in the seed instructions, thus maintaining a smoother difficulty gradient.
The effectiveness of this gradient is further substantiated by the theory presented in Kung et al. (2023)’s work. They categorize samples with high instruction uncertainty and a high prediction probability as “ambiguous”. Instruction uncertainty denotes the degree to which minor modifications in the instruction impact response generation. Higher uncertainty indicates greater sensitivity of LLMs to the specific instructions in a sample. Their research underscores the significant role of “ambiguous” training data in fine-tuning LLMs. Conversely, samples characterized by low instruction uncertainty and prediction probability are deemed challenging, potentially offering limited benefits for fine-tuning due to their complexity.
In our study, we employ the methodology of Kung et al. (2023) to evaluate the "ambiguity" of datasets. We simplify the process by calculating instruction uncertainty as the average loss changes upon altering instructions and representing prediction probability by the inverse of response loss. The resulting plot is analogous to that in Kung et al. (2023)’s study.
Figure 4 exemplifies this concept with the model fine-tuned using $\textbf{C}_{30}$. Here, red spots denote the ambiguity of $\textbf{H}_{20}$, while blue points represent $\textbf{C}_{30r}$. The data reveals that instructions from $\textbf{C}_{30r}$ are challenging, evidenced by their lower instruction uncertainty and prediction probabilities. In contrast, instructions from $\textbf{H}_{20}$ are more “ambiguous”, indicated by higher prediction probabilities and uncertainties, thus making them more conducive for the fine-tuning process. From these observations, we conclude that the Instruction Fusion (IF) method effectively generates more “ambiguous” samples. LLMs exhibit increased responsiveness to the fused instructions compared to the original seed instructions.
@@
Instruction Fusion (IF) counters the diversity limitations of Evol-Instruct, which is constrained by initial objectives, by creating instructions that blend the targets of parent instructions. This leads to unique tasks not found in the Evol-Instruct dataset. In Figure 5, the gray points represent a t-SNE 2D sentiment embedding plot for $\mathbf{C}_{30}$ instructions, showing small cliques indicative of limited diversity. The orange points represent samples from $\mathbf{C}_{30r}$. We can observe that the additional evolved data doesn’t fill the gap between those small cliques, which can demonstrate how the instruction diversity provided by Evol-Instruct is limited.
Conversely, the red points, representing IF’s fused instructions, populate previously empty areas, demonstrating increased diversity. The red points’ distribution fills gaps left by the blue points, illustrating a more uniform sentiment embedding.
To quantify this dispersion, we calculate the variance of nearest neighbor distances for all instructions:
@@
The results reveal notable findings. The first plot of $\textbf{C}_{\text{PI}}$, composed of $\textbf{C}_{30}+\textbf{C}_{30r}$, exhibits higher dispersion with a variance of 0.0332, indicating reduced uniformity compared to $\textbf{C}_{30}$ alone, which is 0.0316. However, an interesting contrast arises when employing Instruction Fusion on $\textbf{C}_{30}$: by integrating $\textbf{H}_{20}$ with $\textbf{C}_{30}$, the dispersion of the second plot decreases by 28.5% from 0.0332 to 0.0226. This significant reduction suggests that Instruction Fusion effectively narrows the gaps in sentiment embeddings between existing instructions. This improvement is achieved by combining the objectives of two distinct instructions, demonstrating the method’s efficacy in enhancing instruction diversity.
In conclusion, the Instruction Fusion (IF) method we propose serves as a valuable complement to Evol-Instruct, effectively addressing its limitations. By amalgamating diverse objectives from various seed instructions, IF leads to the creation of a more challenging and diverse set of instructions. This method not only expands the range of tasks addressed but also ensures a more harmonious and gradual escalation in the complexity of instructions.
@@ 3 Experiments
IF-CL-EC 34B weight&data 78.7 71.3 71.4 60.7
In this section, we report the experiment details of LLMs fine-tuned with instruction generated by IF. We focused on the pass@1 performance under greedy generation settings across five benchmarks: four Python benchmarks (HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), ${}^{+}$, ${}^{+}$ Liu et al. (2023a)) and the multi-language benchmark MultiPL-E Cassano et al. (2022). The official EvalPlus code [^3] https://github.com/evalplus/evalplus was used for evaluating the Python benchmarks, while bigcode-evaluation-harness Ben Allal et al. (2022) was utilized to evaluate the performance on MultiPL-E.
@@ 3.1 Experiment Details
For the selection of our base models, we have chosen to utilize CodeLlama and CodeLlama-Python to evaluate the effectiveness of our proposed method. It’s important to note that while the recently released open-source model, DeepSeek-Coder DeepSeek (2023), has achieved state-of-the-art performance among other base code LLMs, the specifics of their techniques and data were not accessible at the time of writing this paper. Consequently, we have opted not to include DeepSeek-Coder in our experimental analysis.
@@ 3.2 Evaluation
Table 2 presents the performance of various open-source models on Python benchmarks. When compared to the leading closed-source models, GPT-3.5 Turbo, our models with 13B and 34B parameters demonstrate superior performance on the HumanEval and ${}^{+}$ benchmarks. However, there is a noticeable performance gap on the two MPBB benchmarks. This discrepancy may stem from the MBPP test cases, which often require specific prior knowledge for resolution. For instance, in the real test case task_id:20, the prompt "Write a function to check if a given number is woodball or not." requires the model’s understanding of what constitutes a ’woodball’ number. This type of prior knowledge is either not acquired during the pre-training stage or forgotten during the supervised fine-tuning process, leading to a consistent failure across all IF models for this case.
However, compared with all open-source models, the IF models excel in Python code generation performance, surpassing all baselines. The fused instructions, which offer richer sentiment, increased difficulty, and smoother transitions, allow both the 13B and 34B versions of IF models to achieve new benchmarks, outperforming models of comparable sizes. Notably, the 13B version of IF-CLP-ECP outperforms all 34B baselines in the HumanEval and ${}^{+}$ benchmarks and shows comparable results on MBPP and ${}^{+}$.
@@ 3.3 Ablation Study
Method HumanEval HumanEval+ MBPP MBPP+ IF-CLP-ECP 73.8 69.5 71.7 61.7   -fused inst. 67.7 64.0 66.4 56.4 IF-CL-ECP 75.6 68.9 67.9 58.6   -fused inst. 66.5 61.6 67.2 57.4 IF-CL-EC 74.4 68.3 69.7 59.4   -fused inst. 65.2 61.0 68.4 56.1 IF-CLP-ECP 75.6 69.5 73.7 62.7   -fused inst. 72.0 65.2 72.2 61.9 IF-CL-ECP 76.8 70.7 73.2 62.9   -fused inst. 67.7 62.8 70.9 61.2 IF-CL-EC 78.7 71.3 71.4 60.7   -fused inst. 67.7 62.2 69.9 60.4
@@
Further analysis of Instruction Fusion’s effectiveness involved an ablation study using various combinations of IF and the original evolved instructions. Figure 6 depicts the performance of Code Llama when fine-tuned with different volumes of original instructions. The data indicates that the 13B model achieves optimal performance with 30K evolution samples, suggesting that an increased quantity of evolved instructions does not linearly translate to better performance. Moreover, Luo et al. (2023) observed a performance decrease when extending evolution to the fourth round, implying a tangible upper limit to the effectiveness of Code Evol-Instruct in code generation tasks. However, the integration of IF samples with the original evolved dataset transcend this limitation, resulting in a notable enhancement in model performance. Conversely, the 34B model does not show the same constraints with the current range of evolved samples, possibly due to its larger parameter size.
@@ 5 Conclusion
In this paper, we introduce the Instruction Fusion technique, an advancement of Evol-Instruct specifically tailored for code generation tasks. This technique involves the fusion of two evolved instructions into a single, cohesive prompt. It excels in creating instruction sets that are reasonably complex, yet manageable, facilitating a progressive increase in difficulty which is achieved as objectives can be learned separately. The fused instructions also grants a better diversity of the instruction pool. Notably, Large Language Models (LLMs) fine-tuned with this method have demonstrated superior performance across the top five benchmarks, setting new state-of-the-art records.
@@
The primary limitation of the Instruction Fusion method proposed in our study is its cost. We employ GPT-4 Turbo as the Teacher Language Model (LLM), and the fusion prompt incorporates a higher token count compared to the Evol-Instruction. This results in increased expenses for data collection, which cost around 2,200 USD in total. However, it is worth noting compared with manual annotation. Moreover, such costs are rapidly decreasing, thanks to the swift advancements in LLM technology. Another point of concern is the success rate of the fusion process. Fusion of Python-based instructions achieves a pass rate of 93% (where 7% of the fused instructions are deemed unsolvable by GPT-4 Turbo). In contrast, cross-language fusion exhibits a lower pass rate of approximately 63%. This discrepancy arises because many tasks are typically resolved using specific programming languages. Consequently, fusing instructions from different programming languages often leads to impractical outcomes. To address this issue, we propose the development of new prompts that categorize tasks prior to sampling, a solution which could be to explore in future research.